The product of a number and -6 amounts to five times the sum of that number and 33.find the number
Found 2 solutions by lwsshak3, ikleyn:
Answer by lwsshak3(11628)

The product of a number and -6 amounts to five times the sum of that number and 33.find the number
***
let x=the number
-6x=5x+33
-11x=33
x=-3
the number=-3

Answer by ikleyn(53961)

.
The product of a number and -6 amounts to five times the sum of that number and 33.
find the number
~~~~~~~~~~~~~~~~~~~~~~~~~~~


        I read this problem differently from that by @lwsshak3.
        My solution and my answer are different.


-6x = 5(x+33)
-6x = 5x + 165
-11x = 165
x = 165/(-11) = -15.

ANSWER. The number is -15.